# Statistical thermodynamics of the human brain activity, the Hagedorn temperature and the Zipf law

### v1

@@ Abstract
Summarizing, the analysis of large scale fMRI brain recordings via a machine learning approach shows the existence of a robust region of a linear $S(E)$ regime, whose range and proximity to the “operating temperature” $T=1$ increases with the dimensionality of the sampling $N$, i.e. going to finer and finer length-scales. This leads to the finding that the brain operates asymptotically at the Hagedorn temperature. The appearance of the linear $S(E)$ dependence demonstrates that a wide range of brain states follow Zipf’s law. The approach presented in this paper on one side allowed the first empirical demonstration of the wide linear regime of a very diverse (critical) density of brain states – fundamental for normal brain function – and on the other offers a powerful tool to study other complex systems.

### v2

@@ IV Discussion
Summarizing, the analysis of large scale fMRI brain recordings via a novel machine learning approach shows the existence of a robust region of a linear $S(E)$ regime, whose range and proximity to the “operating temperature” $T_{ref}=1$ increases with the dimensionality of the sampling $N$, i.e. going to finer and finer length-scales. This leads to the finding that the brain operates asymptotically at the Hagedorn temperature. The appearance of the linear $S(E)$ dependence demonstrates that a wide range of brain states follow Zipf’s law. The approach presented in this paper on one side allowed the first empirical demonstration of the wide linear regime of a very diverse (critical) density of brain states – fundamental for normal brain function – and on the other offers a powerful tool to study other complex systems.